OpenStudy (anonymous):

Write an equation of a line in slope-intercept form that is parallel to x + 4y = 6 and passes through the point (-8, 5) A. y=−1 4 x+3 B. y=−1 4 x+6.75 C. y = x - 13 D. y = x + 13

OpenStudy (solomonzelman):

x + 4y = 6 4y=-x+6 y=-1/4x+6 so lets do the following y=-1/4x+b plug in (-8, 5) 5=-1/4(8)+b solve for b and rewrite it.

OpenStudy (anonymous):

what do you mean?

OpenStudy (solomonzelman):

solve for b or for the y intercept.

OpenStudy (solomonzelman):

Do you get how I got y=-1/4x+6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that's the problem I don't know how to get b